OpenStudy (anonymous):

help please!!! (: Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6). 7 −14 0 14

OpenStudy (blurbendy):

Use the formula Y2 - Y1 / X2 - X1 -6 -8 = -14 -3 +4 = 1 -14/ 1 = -14 slope = -14

OpenStudy (anonymous):

thank you so much!(:

OpenStudy (blurbendy):

np!